OpenStudy (anonymous):

What is the general form of the equation of a circle with a center at (–5, 0) and a radius of 4?

OpenStudy (anonymous):

\[(x-h)^2+(y-k)^2=r^2\] in this case \[h=-5, k = 0, r = 4\] you get \[(x+5)^2+y^2=16\]

OpenStudy (anonymous):

don't write \[(y-0)^2\] or your teacher will think you are a moron

OpenStudy (anonymous):

haha thanks tho

OpenStudy (anonymous):

yw